7) Consider the following reaction:
CaCO3(s)→CaO(s)+CO2(g).
Estimate ΔG∘ for this reaction at each of the following temperatures. (Assume that ΔH∘ and ΔS∘ do not change too much within the given temperature range.)
part a) 300 K
part b) 1055 K
part c)1480 K

8) Determine ΔG∘ for the following reaction:

2NO(g)+O2(g)→N2O4(g)

Use the following reactions with known ΔG∘rxn values:

N2O4(g)→2NO2(g), ΔG∘rxn = 2.8 kJ
NO(g)+12O2(g)→NO2(g), ΔG∘rxn = - 36.3 kJ
1
Expert's answer
2020-04-13T02:27:43-0400

7)


"\Delta G _0 = -RTlnK"

Supposably


"K = const"

a)


"\Delta G _0 = -8.314*300lnK"

b)


"\Delta G _0 = -8.314*1055lnK"

The parameter will be lower by 3.5 than in the case of a

с)


"\Delta G _0 = -8.314*1480lnK"

The parameter will be lower by 1.4 than in the case of b and will be lower by 4.9 than in the case of a


8)


"\Delta G_0 = -2.8 +2*(-36.3) = -75.4 kJ"